wiki:Research/LhARA/DesignAndIntegration/Meetings/2023/2023-03-10

Design and Integration meeting: 10 March 2023 16:00 GMT

Zoom: https://ukri.zoom.us/j/92932885675

Agenda

  • Actions from last meeting.
  • Simulations.
  • Preparation of 6-month review document.
  • AOB

Actions

Actions Description Status
22-12-06-04 Chris Baker to provide a CAD model of the Gabor Lens test bench In progress
22-12-06-05 Clive Hill to import the CAD model into Creo to provide a better image of the Gabor Lenses. In progress
23-01-17-02 Ajit Kurup to get Dockerfiles for BDsim from Will. Closed
23-01-17-03 Ajit Kurup to investigate options for storing data somewhere easily accessible for the collaboration. In progress
23-01-24-01 All to prepare document for the milestone to update the baseline In progress

Notes

Present: Neil Bliss, Clive Hill, Ajit Kurup, Ken Long, Hywel Owen, Jaroslaw Pasternak, Will Shields, Colin Whyte
Apologies: Rehanah Razak

  • Simulations, see slides from Will.
    • Tested emittance growth with just a drift.
    • GPT sampling in time and not using screens does not show the emittance growth.
    • No idea why using screens has this issue.
    • HO: Can contact people at LD who use GPT regularly to see if they can help.
    • HO posted a link to this article https://journals.aps.org/prab/abstract/10.1103/PhysRevSTAB.6.034202
    • NB: What are the details of the nozzle?
      • 2.87mm is the exit aperture radius. Can increase this aperture to 4mm to improve transmission.
  • Document preparation.
    • NB circulated the pdf and has sent Hywel link to the Word
  • CAD, slides from Clive.
    • CAD of Gabor lens and target chamber for SCAPA.
    • A tape drive target has been added to the drawing.
    • The diagnostic camera behind the target requires a target to end flange distance of 130mm.
    • Discussion on the different vessels needed for the laser target.
    • HO gave some details of the EPAC target.
    • Discussion of details of the target region.
    • HO: Where does the waste laser go? This is an issue for EPAC.
    • A meeting with Ross and Chris should be arranged to discuss these details.

  • AOB
    • Agreed to cancel the meeting next Tuesday as we will meet on Friday.
Last modified 12 months ago Last modified on Mar 16, 2023 4:11:35 PM

Attachments (1)

Download all attachments as: .zip